OpenStudy (anonymous):

Deterime whether two real numbers whose sum is 17 can have each of the following products. If so, find the real numbers. a) 60 b) 52 c) 80

OpenStudy (savvy):

they can have 60 and 52 as products but not 80....

OpenStudy (anonymous):

how did you find that?

OpenStudy (anonymous):

the highest possible product you can get is (17/2)^2

OpenStudy (anonymous):

a) x(17-x) = 60 -x^2+17x=60 -(x-8,5)^2 + 72.25=60 (x-8,5)^2 = 12.25 x-8,5 = sqrt(12.25) x = 5

OpenStudy (anonymous):

what is x?

OpenStudy (anonymous):

x is one of the 2 numbers. the product I wrote on the first line was x (on of the numbers) times 17 -x (since the sum of both numbers is 17, 17 minus one of the numbers will give you the other one)

OpenStudy (anonymous):

for b) it's 4 and 13, you can use the same strategy

OpenStudy (anonymous):

thank u soo much :D

OpenStudy (anonymous):

Nicely done

OpenStudy (savvy):

let the numbers be x and y.... now, x+y = 17 xy=P P:product of x and y now y=17-x so x(17-x)=P 17x - x^2 = P x^2 - 17x + P = 0 it's discriminent, D = 289 - 4P only for P=60 and P=52 D>0 which means only for these values there is a real value of x.

OpenStudy (anonymous):

@ m_charron2 I don't get the part for the second number a)

OpenStudy (anonymous):

I skipped the y step. So you're looking for 2 numbers, let's take one number as x, the other as y As Savvy did, x+y =17 and x*y = 60 so, you isolate one (I chose to isolate y) : y =17 - x Then substitue y in the second equation : x*(17-x) = 60 Once you get x, simply replace it in either equation to get y

OpenStudy (anonymous):

ok thnxx